OpenStudy (anonymous):

4x + 2y = 12, solve for y

OpenStudy (aaronandyson):

4x + 2y = 12 2y = 12 - 4x \[\frac{ 12 - 4x }{ 2 } = y\]

OpenStudy (anonymous):

y=4

OpenStudy (anonymous):

4x + 2y = 12 -4x -4x 2y=-4x+12 2y/2=-4x/2+12/2 y=-2x+6